# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v2

@@ 3.2.1 MARS-AdamW
The first instance of MARS is built up on the idea of Adam/AdamW (Loshchilov and Hutter, 2019). To automatically adjust the learning rate and accelerate convergence, Adam (Kingma and Ba, 2015) adopts the adaptive preconditioned gradient in (3.6) together with a bias correction and $\ell_{2}$ regularization. AdamW (Loshchilov and Hutter, 2019) further changes the $\ell_{2}$ regularization to a decoupled weight decay. Overall, the full AdamW updates can be summarized as follows:
@@
$\mathbf{c}_{t}$ $:=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma_{t}}\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)},$ (3.12)
@@
Here, $\gamma_{t}$ is a scaling parameter that controls the strength of gradient correction. When $\gamma_{t}=0$, the algorithm reduces to AdamW. Conversely, when $\gamma_{t}=1$, (3.13) aligns with the STORM momentum. Combining (3.12), (3.13), (3.14) together with (3.11) and the mirror descent update (3.2), we derive the MARS-AdamW algorithm in Algorithm 2. In practice, $\gamma_{t}$ is often set between 00 and $1$. Moreover, we employ gradient clipping-by-norm to $\mathbf{c}_{t}$ at Line 5, following the standard gradient clipping technique performed in neural network training. We provide a convergence analysis of Algorithm 2 in Theorem 4.6 in Appendix 4.
@@
Connection between MARS-AdamW and Adan.
Adan (Xie et al., 2024) is another adaptive gradient methods improved upon Adam with reformulated Nesterov’s accelerated SGD (See Lemma 1 in Xie et al. (2024) for more details). The Adan algorithm takes the following momentum updates:
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\beta_{1}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\big{]},$
@@ 3.2.2 MARS-Lion
Setting $\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{1}$, $a_{2}=1-\beta_{1}$, $b_{1}=\beta_{2}$, $b_{2}=1-\beta_{2}$ in Lemma C.1, and shifting the index of $\mathbf{u}_{t}$ by taking $\mathbf{u}_{t}=\mathbf{u}_{t-1}$, $\mathbf{u}_{t-1}=\mathbf{u}_{t-2}$ in Lemma C.1, we can show that Lion momentum updates in (3.15) and (3.16) are equivalent to the following single momentum update:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\Big{]}.$ (3.19)
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]}.$ (3.20)
@@
Connection between MARS-Lion and Lion.
Lion turns out to be a special case of MARS-Lion. The momentum updates in Lion can be seen as an approximate implementation of our updates. To facilitate this claim, we present a lemma that follows directly from straightforward arithmetic calculations.
@@ 3.2.3 MARS-Shampoo
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\mathbf{L}_{T}^{-1/2}\mathbf{m}_{t}\mathbf{R}_{T}^{-1/2},$ (3.21)
@@
To reduce the time complexity of SVD decomposition, Bernstein and Newhouse (2024) summarized four different approaches for computing (3.22) including SVD, sketching (Martinsson and Tropp, 2020), Newton iteration (Lakić, 1998; Higham, 2008; Anil et al., 2020), and Newton-Schulz iteration (Schulz, 1933; Higham, 2008). Our algorithm design accommodates any of these SVD solvers to best fit specific computational needs.
Connection between MARS-Shampoo and Muon.
Muon (Jordan et al., 2024) is a recently proposed algorithm that utilizes the Newton-Schulz iteration (Higham, 2008; Schulz, 1933) to solve the SVD problem. It has demonstrated superior performance in terms of convergence speed when compared with AdamW and Shampoo in training large language models. The update rules of Muon are demonstrated as follows:
@@
Applying Lemma C.1 to (3.23) and (3.24), with $\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\mu$, $a_{2}=1$, $b_{1}=\mu$, $b_{2}=1$, we obtain an equivalent single update of momentum:
$\mathbf{m}_{t}$ $=\mu\mathbf{m}_{t-1}+\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\mu\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}.$ (3.25)
@@
$\mathbf{m}_{t}$ $=\mu\mathbf{m}_{t-1}+(1-\mu)\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\mu(1-\mu)\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}.$
@@
$\frac{\mathbf{m}_{t}}{1-\mu}$ $=\mu\cdot\frac{\mathbf{m}_{t-1}}{1-\mu}+\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\mu\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}.$ (3.26)
@@
3.3 Nesterov’s Acceleration is A Speical Case of MARS
Many optimization algorithms exhibit similarities with Nesterov’s acceleration and can be considered as adaptations of Nesterov’s acceleration with varying parameterizations, as discussed in works such as Defazio et al. (2024) and Xie et al. (2024), etc. In this subsection, we compare and contrast Nesterov’s momentum and STORM momentum used in our paper. Specifically, Nesterov’s acceleration (Nesterov, 1983) follows the update rules:
$\mathbf{v}_{t+1}$ $=\beta_{1}\mathbf{v}_{t}+\nabla f(\bm{\theta}_{t}-\eta\beta_{1}\mathbf{v}_{t},\bm{\xi}_{t})$
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\eta\mathbf{v}_{t+1}.$
Taking $\mathbf{x}_{t}=\bm{\theta}_{t}-\eta\beta_{1}\mathbf{v}_{t}$, we have $\bm{\theta}_{t+1}=\mathbf{x}_{t}-\eta\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$. Nesterov’s acceleration reduces to the following form:
$\mathbf{x}_{t+1}$ $=\bm{\theta}_{t+1}+\beta_{1}(\bm{\theta}_{t+1}-\bm{\theta}_{t})$
$=\mathbf{x}_{t}+\beta_{1}(\mathbf{x}_{t}-\mathbf{x}_{t-1})-\eta\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\eta\beta_{1}(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1}).$
Defining $\widetilde{\mathbf{m}}_{t}=-\frac{\mathbf{x}_{t+1}-\mathbf{x}_{t}}{\eta}$, we obtain
$\widetilde{\mathbf{m}}_{t}$ $=\beta_{1}\widetilde{\mathbf{m}}_{t-1}+\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\beta_{1}(\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1}))$ (3.27)
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta\widetilde{\mathbf{m}}_{t}.$
On the contrary, the STORM momentum with scaling parameter $\gamma$ is
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\gamma\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]}$
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta\mathbf{m}_{t}.$ (3.28)
One way of comparing (3.27) with (3.28) is to multiply both sides of (3.27) by $(1-\beta_{1})$ and obtain
$(1-\beta_{1})\widetilde{\mathbf{m}}_{t}$ $=\beta_{1}(1-\beta_{1})\widetilde{\mathbf{m}}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\beta_{1}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\Big{]}.$
By defining $\mathbf{m}_{t}:=(1-\beta_{1})\widetilde{\mathbf{m}}_{t}$, Nesterov’s acceleration is equivalently written as:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})+\beta_{1}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\Big{)}$
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\frac{\eta}{1-\beta_{1}}\mathbf{m}_{t}.$ (3.29)
By setting $\gamma=1-\beta_{1}$ and $\eta=\frac{\eta}{1-\beta_{1}}$, and replacing the last randomness term $\bm{\xi}_{t}$ with $\bm{\xi}_{t-1}$ in the generalized STORM momentum (3.28), we can recover Nesterov’s acceleration (3.29). Therefore, Nesterov’s acceleration can be seen as a special case of MARS.
Remark 3.4.
We highlight three key differences between Nesterov’s acceleration and MARS: First, in Nesterov’s momentum, the update is based on subtracting $\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ from $\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})$ to determine a direction of lookahead. In contrast, STORM momentum subtracts $\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})$ from $\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ to cancel out the noise introduced by $\bm{\xi}_{t}$. Although similar in form, this leads to a different convergence property. Our theoretical analysis (Section 4) demonstrates that variance-reduced variants of AdamW achieve an improved convergence rate of $\mathcal{O}(T^{-2/3})$. Empirically, we show that variance-reduced momentum outperforms its approximate counterpart (Section D.2). Second, MARS introduces a more general parameterization framework that not only includes Nesterov’s acceleration but also encompasses AdamW, Lion, Adan, and Muon through appropriate choices of $\gamma$, as we demonstrate in Section 3.2.1, 3.2.2, and 3.2.3, respectively. Finally, MARS applies an additional gradient clipping step to the new gradient estimator $\mathbf{c}_{t}$, which is crucial for practical performance. In contrast, standard Nesterov’s acceleration only applies gradient clipping to the gradient itself.
4 Theoretical Analysis
Although Kingma and Ba (2015) provided a convergence analysis for Adam, Reddi et al. (2019a) pointed out that the analysis is flawed in the convex optimization setting, and they also proved that in some special cases, Adam does not converge. However, there are some attempts to prove the convergence of Adam and AdamW in the nonconvex optimization setting (Zhang et al., 2022b; Li et al., 2024; Zhou et al., 2024). Our proposed algorithm, MARS, is also built upon AdamW and further leverages variance reduction techniques. With a careful choice of hyperparameters, we show that MARS can achieve a faster convergence rate to the first-order stationary points in the nonconvex optimization setting.
To facilitate the analysis of convergence for MARS, we make the following assumptions:
Assumption 4.1 (Bounded Variance).
There exists a positive $\sigma$, such that for any noise $\bm{\xi}$, parameter $\mathbf{x}$, and $\nabla F(\mathbf{x})=\mathbb{E}[\nabla f(\mathbf{x},\bm{\xi})]$, we have:
$\mathbb{E}\big{[}\|\nabla f(\mathbf{x},\bm{\xi})-\nabla F(\mathbf{x})\|_{2}^{2}\big{]}\leq\sigma^{2}.$ (4.1)
Assumption 4.2 ($L$-Smoothness).
For any $\bm{\xi}$, $f(\mathbf{x},\bm{\xi})$ is $L$-smooth:
$\|\nabla f(\mathbf{x},\bm{\xi})-\nabla f(\mathbf{y},\bm{\xi})\|_{2}\leq L\|\mathbf{x}-\mathbf{y}\|_{2},~{}\forall\mathbf{x},\mathbf{y}.$ (4.2)
Assumption 4.3 (Positive Hessian).
There exists a constant $\rho>0$ such that for all $\mathbf{H}_{t},t>0$, $\mathbf{H}_{t}\succ\rho\bm{I}$.
While Assumptions 4.1, 4.2 are commonly used, Assumption 4.3 holds in practice since $\mathbf{H}_{t}$ is typically chosen so that its smallest eigenvalue is greater than a small $\epsilon$ to ensure numerical stability.
Based on the above assumptions, we present Theorem 4.4 for our main algorithm (Algorithm 1), and Theorem 4.6 for MARS-AdamW (Algorithm 2), where an additional weight decay is involved.
Theorem 4.4.
Under Assumptions 4.14.2 and 4.3, when choosing $\eta_{t}=(s+t)^{-1/3},s\geq 1$, $\eta_{t}\leq\rho\cdot(2L)^{-1}$, and suppose $c\geq 32L^{2}\rho^{-2}+1$, $\beta_{1,t+1}=1-c\eta_{t}^{2}$ and $\beta_{2,t+1}=1-\eta_{t}^{6}$, then for any $T\geq s$, the output of Algorithm 1 satisfies
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}$ $\leq\Big{(}2\rho G+\frac{\rho c^{2}\sigma^{2}}{4L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-\frac{\rho^{2}\sum_{t=1}^{T}M_{t+1}}{16L^{2}T^{1/3}}},$
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}^{2}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\Big{(}\frac{16G}{3\rho}+\frac{2c^{2}\sigma^{2}}{3\rho L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-\frac{\sum_{t=1}^{T}M_{t+1}}{6L^{2}T^{1/3}}},$
where $G=F(\mathbf{x}_{1})-\min_{\mathbf{x}}F(\mathbf{x})+\frac{\rho s^{1/3}\sigma^{2}}{16L^{2}}$ and $M_{t+1}$ is defined in (B.2).
Notably, the parameters $\rho,G,c,L$, and $\sigma$ are constants. Therefore, Theorem 4.4 establishes an convergence rate of $\mathcal{O}(\log(T)/T^{2/3})$ to the first-order stationary points, which is faster than that of Adam.  [^2] Some papers (Cutkosky and Orabona, 2019; Huang et al., 2021) use the $\ell_{2}$ norm $\|\cdot\|_{2}$ rather than squared $\ell_{2}$ norm $\|\cdot\|_{2}^{2}$ for characterizing convergence to first-order stationary points. In this case, if we define $\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|+\frac{1}{\eta_{t}}\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|$, then by Jensen’s inequality, the convergence rate would be $\mathcal{O}(\log(T)/T^{1/3})$. .
Remark 4.5.
In Theorem 4.4, $M_{t+1}$ is in the form of $\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}-\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\Big{(}\beta_{1,t+1}(1-\gamma_{t+1})-A_{t+1}\Big{)}^{2}$, where $\bm{\Delta}_{t}$ and $A_{t+1}$ are defined in (B.2) and (B.3), respectively. When $\gamma_{t+1}=1$, we obtain $M_{t+1}=0$, which recovers the same convergence result as SuperAdam (Huang et al., 2021). Notably, there exists an optimal sequence of $\gamma_{t}$, i.e., $\gamma_{t+1}=1-\frac{A_{t+1}}{\beta_{1,t+1}}$, such that $M_{t+1}=\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}$, leading to a tighter bound in Theorem 4.4. While the dominant convergence rate remains the same, this refined bound consistently yields a smaller upper bound due to the non-positive term (highlighted in red color), leading to improved practical performance.
Theorem 4.6.
Under Assumptions 4.14.2 and 4.3, when choosing $\eta_{t}=(s+t)^{-1/3},s\geq 1$, $\eta_{t}\leq\min\{(4\lambda)^{-1},\rho\cdot(2L)^{-1}\}$., and suppose $c\geq 32L^{2}\rho^{-2}+1$, $\beta_{1,t+1}=1-c\eta_{t}^{2}$ and $\beta_{2,t+1}=1-\eta_{t}^{6}$. In addition, assume that there exists a constant $D>0$ such that $\|\mathbf{x}_{t}\|_{2}\leq D$ for all $t>0$. Then for any $T\geq s$, the output of Algorithm 2 satisfies
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}$ $\leq\Big{(}2\rho(G+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda D^{2}\log(s+T)})+\frac{\rho c^{2}\sigma^{2}}{4L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-\frac{\rho^{2}\sum_{t=1}^{T}M_{t+1}}{16L^{2}T^{1/3}}},$
$\frac{1}{T}\sum_{t=1}^{T}\frac{1}{\eta_{t}^{2}}\cdot\mathbb{E}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}$ $\leq\Big{(}\frac{16(G+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda D^{2}\log(s+T)})}{\rho}+\frac{2c^{2}\sigma^{2}}{\rho L^{2}}\cdot\log(s+T)\Big{)}\cdot\frac{1}{T^{2/3}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-\frac{\sum_{t=1}^{T}M_{t+1}}{2L^{2}T^{1/3}}}.$
where $G=F(\mathbf{x}_{1})-\min_{\mathbf{x}}F(\mathbf{x})+\frac{\lambda}{2}D^{2}(1+\epsilon)+\frac{\rho s^{1/3}\sigma^{2}}{16L^{2}}$ and $M_{t+1}$ is defined in (B.2).
Theorem 4.6 guarantees an $O(\log(T)/T^{2/3})$ convergence rate to the first-order stationary points for MARS-AdamW, which is faster than that of AdamW. Notably, the involvement of weight decay introduces an additional term $\lambda D^{2}\log(s+T)$ (highlighted in blue color), yet the overall convergence rate remains unchanged. Note that the weight decay has been shown to improve the generalization performance of the optimizer, à la Loshchilov and Hutter (2019). Finally, similar to Theorem 4.4, the term $M_{t+1}$ is always non-negative, and when $\gamma_{t+1}=1$, we obtain $M_{t+1}=0$. This indicates that by adopting a flexible $\gamma_{t}$ schedule, the upper bound can always be made tighter due to the non-positive term (highlighted in red color), leading to improved theoretical and empirical performance.

### v3

@@ 3.2.1 MARS-AdamW
The first instance of MARS is built up on the idea of Adam/AdamW (Loshchilov & Hutter, 2019). To automatically adjust the learning rate and accelerate convergence, Adam (Kingma & Ba, 2015) adopts the adaptive preconditioned gradient in (3.6) together with a bias correction and $\ell_{2}$ regularization. AdamW (Loshchilov & Hutter, 2019) further changes the $\ell_{2}$ regularization to a decoupled weight decay. Overall, the full AdamW updates can be summarized as follows:
@@
$\mathbf{c}_{t}$ $:=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$+{\color[rgb]{1,0,0}\gamma_{t}}\frac{\beta_{1}}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)},$ (3.12)
@@
Here, $\gamma_{t}$ is a scaling parameter that controls the strength of gradient correction. When $\gamma_{t}=0$, the algorithm reduces to AdamW. Conversely, when $\gamma_{t}=1$, (3.13) aligns with the STORM momentum. Combining (3.12), (3.13), (3.14) together with (3.11) and the mirror descent update (3.2), we derive the MARS-AdamW algorithm in Algorithm 2. In practice, $\gamma_{t}$ is often set between 00 and $1$. Moreover, we employ gradient clipping-by-norm to $\mathbf{c}_{t}$ at Line 5, following the standard gradient clipping technique performed in neural network training. We provide a convergence analysis of Algorithm 2 in Theorem B.6 in Appendix B.2.
@@
Connection between MARS-AdamW and Adan. Adan (Xie et al., 2024) is another adaptive gradient method improved upon Adam with reformulated Nesterov’s accelerated SGD (See Lemma 1 in Xie et al. (2024) for more details). The Adan algorithm takes the following momentum updates:
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$\qquad+\beta_{1}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\big{]},$
@@ 3.2.2 MARS-Lion
Setting $\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{1}$, $a_{2}=1-\beta_{1}$, $b_{1}=\beta_{2}$, $b_{2}=1-\beta_{2}$ in Lemma D.1, and shifting the index of $\mathbf{u}_{t}$ by taking $\mathbf{u}_{t}=\mathbf{u}_{t-1}$, $\mathbf{u}_{t-1}=\mathbf{u}_{t-2}$ in Lemma D.1, we can show that Lion momentum updates in (3.15) and (3.16) are equivalent to the following single momentum update:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.19)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\Big{]}.$ (3.20)
@@
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.21)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]}.$ (3.22)
@@
Connection between MARS-Lion and Lion. Lion turns out to be a special case of MARS-Lion. The momentum updates in Lion can be seen as an approximate implementation of our updates. To facilitate this claim, we present a lemma that follows directly from straightforward arithmetic calculations.
Lemma 3.4.
For any sequence $\{\mathbf{g}_{t}\in\mathbb{R}^{d}\}_{t=0,1,\ldots}$, consider the following updates of $\mathbf{m}_{t}$ for any constant factors $a_{1},a_{2},b_{1}$, and $b_{2}$:
$\mathbf{u}_{t}$ $=a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{g}_{t},$ (3.23)
$\mathbf{m}_{t}$ $=b_{1}\mathbf{u}_{t}+b_{2}\mathbf{g}_{t}.$ (3.24)
The updates are equivalent to
$\mathbf{m}_{t}=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{g}_{t}+a_{1}b_{2}(\mathbf{g}_{t}-\mathbf{g}_{t-1}).$
Setting $\mathbf{g}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{1}$, $a_{2}=1-\beta_{1}$, $b_{1}=\beta_{2}$, $b_{2}=1-\beta_{2}$ in Lemma D.1, and shifting the index of $\mathbf{u}_{t}$ by taking $\mathbf{u}_{t}=\mathbf{u}_{t-1}$, $\mathbf{u}_{t-1}=\mathbf{u}_{t-2}$ in Lemma D.1, we can show that Lion momentum updates in (3.15) and (3.16) are equivalent to the following single momentum update:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.25)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\Big{]}.$ (3.26)
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.27)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]}.$ (3.28)
@@ 3.2.3 MARS-Shampoo
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\mathbf{L}_{t}^{-1/4}\mathbf{m}_{t}\mathbf{R}_{t}^{-1/4},$ (3.29)
@@
To reduce the time complexity of SVD decomposition, Bernstein & Newhouse (2024) summarized four different approaches for computing or approximating (3.30) including SVD, sketching (Martinsson & Tropp, 2020), Newton iteration (Lakić, 1998; Higham, 2008; Anil et al., 2020), and Newton-Schulz iteration (Schulz, 1933; Higham, 2008). Our algorithm design accommodates any of these SVD solvers to best fit specific computational needs.
Connection between MARS-Shampoo and Muon. Muon (Jordan et al., 2024) is a recently proposed algorithm that utilizes the Newton-Schulz iteration (Higham, 2008; Schulz, 1933) to solve the SVD problem. It has demonstrated superior performance in terms of convergence speed when compared with AdamW and Shampoo in training large language models. The update rules of Muon are demonstrated as follows:
@@
Applying Lemma D.1 to (3.31) and (3.32), with $\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\mu$, $a_{2}=1$, $b_{1}=\mu$, $b_{2}=1$, we obtain an equivalent single update of momentum:
$\mathbf{m}_{t}$ $=\mu\mathbf{m}_{t-1}+\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$\qquad+\mu\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}.$ (3.33)
@@
$\mathbf{m}_{t}$ $=\mu\mathbf{m}_{t-1}+(1-\mu)\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$\qquad+\mu(1-\mu)\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}.$
@@
$\frac{\mathbf{m}_{t}}{1-\mu}$ $=\mu\cdot\frac{\mathbf{m}_{t-1}}{1-\mu}+\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$\qquad+\mu\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}.$ (3.34)